OpenStudy (anonymous):

In 2002, hourly wages for the top three CEOs of U.S corporations were calculated based on their working 14 hours per day for 365 days. Together the three CEOs made $44,000 per hour. The top CEO earned $2,000 more per hour than the combined hourly wages of the next two top CEOs, and the top CEO earned $7,000 more per hour than the next top CEO. (Source: USA today) (A) let x, y, and z be the hourly wages in thousands of dollars for the top CEOs from greatest to least. Write a system of equations whose solution gives these hourly wages. (B) solve the system. Interpret the answer.

OpenStudy (jdoe0001):

so we have 3 CEO's hourly wages Mr x, Mr y and Mr z their wages combined are $44,000 so \(\bf \color{blue}{x + y + z = 4000}\) we also know that the top CEO, Mr x, makes whatever the other 2 guys make PLUS $2,000 so, how much do the other guys make? Mr y and Mr z, well, they make (y + z) thus Mr x makes \(\bf \color{blue}{(y+z)+2000 = x}\) we also know that the top CEO Mr x, makes more than Mr y the next below in the ladder, so how much? well Mr x makes whatever Mr y makes PLUS $7,000 so \(\bf \color{blue}{y + 7000 = x}\)

OpenStudy (anonymous):

So what steps do I have to do first to solve this?